OpenStudy (anonymous):

One jug holds 7/8 gallon of water. If a faucet can release 16 gallons of water in 8 minutes, how long would it take to fill one jug with water

OpenStudy (mathstudent55):

This is a proportion problem: 16 gal : 8 minutes = 7/8 gal : x minutes 16/8 = (7/8)/x Cross multiply and solve for x